In the accompanying table, you are given information about a firm that operates in a competitive...

  1. In the accompanying table, you are given information about a firm that operates in a competitive market. Assume that TFC is $20.

Q

TVC

MC

AVC

1

$24

2

30

3

38

4

48

5

62

6

82

7

110

  1. Complete the table (solve for MC and AVC). Show sample calculations.

  1. What is the lowest P at which the firm will produce in the short run?
  2. How many units of output will it produce at that P (referring to b)?
  3. How many units will it produce if the market price is $20?
  4. At P=$20, is this a profit or loss situation for the firm? Calculate the profit or loss.
Answer #1

a.

TFC TVC MC $24 30 38 AVC $24 Q 24 1 20 $6 $15 2 20 $13 $8 3 20 $10 $12 4 48 20 $14 $12 62 20 $20 $14 82 20 $28 $16 110 20 CO

MC=TCn-TCn-1

MC of second unit of output=30-24

=$6

AVC=TVC/Q

AVC of second unit of output=30/2

=$15

b.

Since shut-down condition is

P=MC=minimum of AVC

Since MC cuts AVC at its minimum point. So when MC is greater than AVC, then it means it is the price at which firm will continue to produce in the short-run because. It means minimum price is $14.

c.

The profit-maximizing condition of the perfectly competitive firm in the short-run will be

P=MC

Hence corresponding to this condition firm will produce 5 units of output.

d.

If price=$20

firm will produce 6 units of output because firm MC and P are equal.

e.

At Q=6

TVC=82

TFC=20

TC=82+20

=102

TR=P*Q

=20*6

=120

Profit=TR-TC

=120-102

=$18